# Assessment of Simulation-based Inference Methods for Stochastic Compartmental Models

### v2

@@ Abstract
Global pandemics, such as the recent COVID-19 crisis, highlight the need for stochastic epidemic models that can capture the randomness inherent in the spread of disease. Such models must be accompanied by methods for estimating parameters in order to generate fast nowcasts and short-term forecasts that can inform public health decisions. This paper presents a comparison of two advanced Bayesian inference methods: 1) pseudo-marginal particle Markov chain Monte Carlo, short Particle Filters (PF), and 2) Conditional Normalizing Flows (CNF). We investigate their performance on two commonly used compartmental models: a classical Susceptible-Infected-Recovered (SIR) model and a two-variant Susceptible-Exposed-Infected-Recovered (SEIR) model, complemented by an observation model that maps latent trajectories to empirical data. Addressing the challenges of intractable likelihoods for parameter inference in stochastic settings, our analysis highlights how these likelihood-free methods provide accurate and robust inference capabilities. The results of our simulation study further underscore the effectiveness of these approaches in capturing the stochastic dynamics of epidemics, providing prediction capabilities for the control of epidemic outbreaks. Results on an Ethiopian cohort study demonstrate operational robustness under real‑world noise and irregular data sampling. To facilitate reuse and to enable building pipelines that ultimately contribute to better informed decision making in public health, we make code and synthetic datasets publicly available.
@@ 1 Introduction
The COVID-19 pandemic underscored the critical importance of epidemic modeling research. Rapid, data-driven approaches have demonstrated their capacity to provide public health decision-makers with essential insights for mitigating emerging infectious disease outbreaks and controlling epidemic dynamics [ethiopian_data, doms2018assessing]. Beyond COVID-19, other outbreaks such as Ebola and MPOX have likewise been designated as “public health emergencies of international concern” by the World Health Organization (WHO) [gpmb2024]. These crises and consequent public health challenges highlight the need for advanced modeling techniques capable of capturing the intricate dynamics of disease spread.
@@
Epidemiological and medical research revealed that the spread of infectious diseases is shaped by a variety of processes, including transmission dynamics, patient responses to treatment, and pathogen mutation. A wide spectrum of mathematical modeling approaches is used to describe the underlying dynamics, spanning from standard ordinary differential equation (ODE) models [McKendrick_Kermack, hethcote2000] to agent-based models [bruch2015agent]. On the one hand, ODE-based population models provide a fast to simulate framework, but fall short to describe the variability of events and the resulting complex, stochastic interactions, due to their deterministic nature. Therefore, stochastic models are essential for realistically representing infectious disease dynamics observed in real-world settings, especially in small populations or during the early stages of an outbreak [britton2010stochastic]. On the other hand, purely stochastic models, such as agent-based and continuous-time Markov chain (CTMC) models [dragatz_book, infect_dis_modelling] provide highly detailed stochastic representations of individual-level processes, but come at the cost of prohibitive computational burden [petzold_2013]. Mathematical models based on stochastic differential equations – so called stochastic (meta-)population models – can solve the trade-off between fast model simulations and keeping the stochasticity in the model [dragatz_paper]. In this study, we consider compartmental models described by stochastic differential equations (SDE). In compartmental models individuals are arranged in a finite number of mutually disjoint compartments and interaction happens through transfer from one compartment to another which is described by transition equations [Brauer2008]. Their description in terms of SDEs can be seen as a mesoscopic view on the model dynamics. The SDEs are derived as a diffusion approximation to the microscopic, individual-based, discrete-state CTMC description [dragatz_paper, expansion_vanKampen, fearnhead2014inference] and in the large population limit resemble the macroscopic, population-based, (deterministic) ODE models [kurtz1970solutions]. For a more detailed discussion on stochastic epidemic modeling see [infect_dis_modelling, dragatz_paper, AnderssonAndBritton] and references contained therein.
@@
Many methods have been developed to address the issue of intractable likelihoods and posterior distributions. Early attempts used data augmentation or employed Markov chain Monte Carlo (MCMC) methods [dragatz_book, gibson1998estimating, roberts1999bayesian]. Later approaches sought to reduce computational cost through surrogate models such as Gaussian processes [swallow2022challenges] or linear noise approximations of the underlying Markov jump process [fintzi2022linear]. Hybrid strategies combining data augmentation with MCMC have also been proposed [mckinley2014simulation, el2016bayesian]. Sequential Monte Carlo methods were introduced as a flexible approach for analyzing stochastic state space models and performing Bayesian filtering tasks [doucet_sequential, chopin2013]. These methods, also known as particle filters (PF), have been used to generate parameter proposals of an MCMC algorithm, leading to pseudo-marginal methods for Bayesian inference on model parameters [andrieu_roberts, andrieu_2010, chopin_2020], a popular method for exact Bayesian inference in stochastic models [chopin_2020, wiqvist_2021].
@@
Instead of determining model parameters via sampling based on evaluations of the likelihood or an approximation to it, it is possible to conduct the inference procedure solely based on simulations. In the widely used technique of Approximate Bayesian Computing (ABC), simulated and observed data is compared based on summary statistics [rubin1984, minter_2019]. However, manually selected summary statistics can be uninformative and their selection poses a challenge on its own. Additionally, as these methods require massive numbers of simulations, more recent machine learning approaches have been adapted for simulation-based inference with neural networks [cranmer2020frontier], namely neural posterior estimation methods. These promising techniques try to circumvent summary statistics and directly approximate the distribution of the parameters from samples generated by the simulator. Therein, conditional normalizing flows (CNF) are an advanced method that uses invertible neural networks to generate samples from the posterior distribution of the parameters [radev2020bayesflow].
@@
PF and CNFs show great potential for Bayesian inference on stochastic models and are increasingly recognized within the epidemiological modeling community for providing probabilistic estimates of disease characteristics to effectively monitor pandemic outbreaks [radev2021outbreakflow, storvik2023]. However, their adoption has been limited by the lack of a comprehensive comparison assessing their relative performance. This work offers a praxis-driven comparison between both inference methods, assessing strengths and weaknesses of the particular approaches, as well as their alignment. In the first part, we described the two stochastic compartmental models, the well-known SIR model [McKendrick_Kermack] and a two-variant SEIR model, which was used to analyze the spread of COVID-19 in Ethiopia [ethiopian_data]. Together with the observational model bridging compartment trajectories with empirical data they serve as the foundation for our study. Moreover, we outline Bayesian inference for stochastic differential equation-based models, and explain the chosen techniques of CNF and PF together with the implementation of our comparison workflow. In the second part of the work, we outline the results from comparing the two inference methods on synthetically generated data for the two chosen compartmental models and additionally using real-data for the two-variant SEIR model. Both methods provide robust and reliable inference results on the stochastic version of the SIR model with synthetically generated data, validating their implementation at hand. Analyzing the agreement of the estimated marginal posteriors and model fits based on different synthetic datasets for the two-variant SEIR model, including (partially) missing data, shows that both methods yield good fits to the data at hand. However, ill-conditioning of the model leads to differences in the shape of the marginal posteriors due to difference in parameter space exploration by the inference methods. Applying both methods to an reparametrized version with parameter dimension reduction can overcome these issues and improves posterior alignment. Inference based on the data from [ethiopian_data] shows the applicability of both methods to real-world data. A discussion of the results, their limitations and further research directions follows in Section 6.
@@ 2 Stochastic Compartmental Models
Compartmental models are a mathematical framework to describe how individuals in a population move between different states and are fundamental for modeling infectious diseases [Brauer2008]. Such a model is characterized by the vector of compartment sizes $X\in\mathbb{R}^{d_{X}}$ and the transitions of individuals between compartments. A transition from compartment $i$ to compartment $j$ can be described in terms of the stochiometric coefficient $s_{i,j}\in\mathbb{R}^{d_{x}}$, the vector of changes to the compartment sizes, and the rate at which the transition happens
$a_{i,j}^{N}(X)=cX_{i}\sum_{k\in I_{i,j}}\frac{X_{k}}{N},$
where $c$ denotes a positive parameter, $I_{i,j}$ is an index set containing all indices $k\neq i$ of compartments involved in that transition, and $N$ gives the total number of individuals in the population. Let $\mathcal{T}$ denote the set of all tuples $(i,j)$, such that a transition from $X_{i}$ to $X_{j}$ exists. Then the time evolution of the compartment sizes can be modeled using a multivariate stochastic differential equation [dragatz_paper]:
@@
$=\sum_{(i,j)\in\mathcal{T}}a_{i,j}(X(t))s_{i,j}dt+\sum_{(i,j)\in\mathcal{T}}\sqrt{a_{i,j}(X(t)))}s_{i,j}s_{i,j}^{T}d\mathbf{B}(t)$
where $\mathbf{B}$ denotes a vector of independent Brownian motions. The representation of compartmental models in terms of multivariate SDEs can capture the randomness of disease spread in their diffusion part $\sigma(X(t))$ and retain a drift component $\mu(X(t))$ for representing the deterministic trend of the population [AnderssonAndBritton, dragatz_book]. This makes them a favorable choice for modeling disease outbreaks and dynamics.
In this work, we consider two examples of SDE-based compartmental models, a simple standard Susceptible-Infected-Recovered (SIR) model [McKendrick_Kermack] and an extended multi-variant Susceptible-Exposed-Infected-Recovered (SEIR) model [hethcote2000] considering two variants of a single pathogen.
Figure 1: Schematic representations of the models and methods. (a) Graph of compartments and possible transitions with corresponding rate parameters for the SIR model. (b) Graph of compartments and possible transitions with corresponding rate parameters for the two‑variant SEIR model. (c) Workflow for the assessment of the Bayesian inference.
@@ 2.1 SIR Model
The standard SIR model (Figure 1(a)) considers three compartments: $S$ (Susceptible), $I$ (Infectious) and $R$ (Recovered). The model describes the number of individuals in these compartments given the transition parameters $\beta$ (the transmission rate) and $\gamma$ (the recovery rate). Due to its simplicity, the solution of the SIR model and its properties can be analyzed analytically [harko2014exact] and it often serves as an initial model for the modeling of unknown diseases [daley1999epidemic]. For consistency across models, in the following we compare recovery in terms of the mean infectious period, $\gamma^{-1}$, rather than $\gamma$. The events dynamics are described by three compartments and the transitions between them (Table 1).
@@ 2.2 Two-Variant SEIR Model
Many extensions of the SIR model have been developed to account for the heterogeneity of various infectious diseases [keeling2008], for example the Susceptible-Exposed-Infectious-Recovered model, in which the Exposed compartment covers the latent period between being infected and becoming infectious. Recently, an extension to the ODE-based SEIR model was introduced to capture the early COVID-19 dynamics in Ethiopia [ethiopian_data]. In addition to a classical SEIR model, this model distinguishes between two virus variants, namely the wild-type and a novel variant, leading to a total of 10 compartments (Figure 1(b)).
The novel variant is more potent in the sense that it shows a prolonged infectious period and that immunity from the wild-type variant does not protect against reinfection with the novel variant. However, an infection with the novel variant type confers immunity against subsequent infections with both the wild-type and the novel variant. The dynamics of the compartment sizes are described by initialization, variant entering and transitions between the compartments (Table 2).
The model exhibits non-identifiability issues in the sense that different parameter combinations, lying along certain curves, can generate nearly indistinguishable epidemic trajectories (see Supplementary Information A.3). Therefore, we also consider a reparametrized version of the model with a reduced number of model parameters to be estimated. We fix $\kappa^{-1}$ and use the following redefined parameters
$r_{0}=\beta\cdot\gamma^{-1},\quad e_{0}=\kappa^{-1}+\gamma^{-1},\quad s_{0}=s\cdot\gamma^{-1},$

### v3

@@ Abstract
Global pandemics, such as the recent COVID-19 crisis, highlight the need for stochastic epidemic models that can capture the randomness inherent in the spread of disease. Such models must be accompanied by methods for estimating parameters in order to generate fast nowcasts and short-term forecasts that can inform public health decisions. This paper presents a comparison of two advanced Bayesian inference methods: 1) pseudo-marginal particle Markov chain Monte Carlo, using an unbiased likelihood estimate obtained by Particle Filter (PF), and 2) Conditional Normalizing Flows (CNF). We investigate their performance on three commonly used compartmental models: A classical Susceptible-Infected-Susceptible (SIS), a Susceptible-Infected-Recovered (SIR) model and a two-variant Susceptible-Exposed-Infected-Recovered (SEIR) model, complemented by an observation model that maps latent trajectories to empirical data. Addressing the challenges of intractable likelihoods for parameter inference in stochastic settings, our analysis highlights how these likelihood-free methods provide accurate and robust inference capabilities. The results of our simulation study further underscore the effectiveness of these approaches in capturing the stochastic dynamics of epidemics, providing prediction capabilities for the control of epidemic outbreaks. Results on an Ethiopian cohort study demonstrate operational robustness under real-world noise and irregular data sampling. To facilitate reuse and to enable building pipelines that ultimately contribute to better informed decision making in public health, we make code and synthetic datasets publicly available.
@@ 1 Introduction
The COVID-19 pandemic underscored the critical importance of epidemic modeling research. Rapid, data-driven approaches have demonstrated their capacity to provide public health decision-makers with essential insights for mitigating emerging infectious disease outbreaks and controlling epidemic dynamics [Gudina et al., 2021; Doms et al., 2018]. Beyond COVID-19, other outbreaks such as Ebola and MPOX have likewise been designated as “public health emergencies of international concern” by the World Health Organization (WHO) [Global Preparedness Monitoring Board, 2025]. These crises and consequent public health challenges highlight the need for advanced modeling techniques capable of capturing the intricate dynamics of disease spread.
@@
Epidemiological and medical research revealed that the spread of infectious diseases is shaped by a variety of processes, including transmission dynamics, patient responses to treatment, and pathogen mutation. A wide spectrum of mathematical modeling approaches is used to describe the underlying dynamics, spanning from standard ordinary differential equation (ODE) models [Kermack et al., 1927; Hethcote, 2000] to agent-based models Bruch and Atwell [2015]. On the one hand, ODE-based population models provide a fast to simulate framework, but fall short to describe the variability of events and the resulting complex, stochastic interactions, due to their deterministic nature. Therefore, stochastic models are essential for realistically representing infectious disease dynamics observed in real-world settings, especially in small populations or during the early stages of an outbreak [Britton, 2010]. On the other hand, purely stochastic models, such as agent-based and continuous-time Markov chain (CTMC) models [Dragatz, 2010; Allen, 2017] provide highly detailed stochastic representations of individual-level processes, but come at the cost of prohibitive computational burden [Gillespie et al., 2013]. Mathematical models based on stochastic differential equations – so called stochastic (meta-)population models – can solve the trade-off between fast model simulations and keeping the stochasticity in the model [Dragatz, 2006]. In this study, we consider compartmental models described by stochastic differential equations (SDE). In compartmental models individuals are arranged in a finite number of mutually disjoint compartments and interaction happens through transfer from one compartment to another which is described by transition equations [Brauer, 2008]. Their description in terms of SDEs can be seen as a mesoscopic view on the model dynamics. The SDEs are derived as a diffusion approximation to the microscopic, individual-based, discrete-state CTMC description [Dragatz, 2006; van Kampen, 1976; Fearnhead et al., 2014] and in the large population limit resemble the macroscopic, population-based, (deterministic) ODE models [Kurtz, 1970]. For a more detailed discussion on stochastic epidemic modeling see Allen [2017]; Dragatz [2006]; Andersson and Britton [2001] and references contained therein.
@@
Many methods have been developed to address the issue of intractable likelihoods and posterior distributions. Early attempts used data augmentation or employed Markov chain Monte Carlo (MCMC) methods [Dragatz, 2010; Gibson and Renshaw, 1998; O’Neill and Roberts, 1999]. Later approaches sought to reduce computational cost through surrogate models such as Gaussian processes [Swallow et al., 2022] or linear noise approximations of the underlying Markov jump process [Fintzi et al., 2022]. Hybrid strategies combining data augmentation with MCMC have also been proposed [McKinley et al., 2014; El Maroufy et al., 2016]. Sequential Monte Carlo methods were introduced as a flexible approach for analyzing stochastic state space models and performing Bayesian filtering tasks [Doucet et al., 2000; Chopin et al., 2012]. These methods, also known as particle filters (PF), have been used to produce an unbiased likelihood estimate for an MCMC algorithm, leading to pseudo-marginal methods for Bayesian inference on model parameters [Andrieu and Roberts, 2009; Andrieu et al., 2010; Chopin and Papaspiliopoulos, 2020], a popular method for exact Bayesian inference in stochastic models [Chopin and Papaspiliopoulos, 2020; Wiqvist et al., 2021].
@@
Instead of determining model parameters via sampling based on evaluations of the likelihood or an approximation to it, it is possible to conduct the inference procedure solely based on simulations. In the widely used technique of Approximate Bayesian Computing (ABC), simulated and observed data is compared based on summary statistics [Rubin, 1984; Minter and Retkute, 2019]. However, manually selected summary statistics can be uninformative and their selection poses a challenge on its own. Additionally, as these methods require massive numbers of simulations, more recent machine learning approaches have been adapted for simulation-based inference with neural networks [Cranmer et al., 2020]. These promising techniques often circumvent hand-crafted summary statistics and directly approximate the distribution of the parameters from samples generated by the simulator. Therein, conditional normalizing flows (CNF) are an advanced method that uses invertible neural networks to generate samples from the posterior distribution of the parameters [Papamakarios and Murray, 2016; Papamakarios et al., 2021].
@@
PF and CNFs show great potential for Bayesian inference on stochastic models and are increasingly recognized within the epidemiological modeling community for providing probabilistic estimates of disease characteristics to effectively monitor pandemic outbreaks [Radev et al., 2021; Storvik et al., 2023]. However, their adoption has been limited by the lack of a comprehensive comparison assessing their relative performance. This work offers a praxis-driven comparison between both inference methods, assessing strengths and weaknesses of the particular approaches, as well as their alignment. In the first part, we described the three stochastic compartmental models, the well-known SIS and SIR models [Kermack et al., 1927] and a two-variant SEIR model, which was used to analyze the spread of COVID-19 in Ethiopia [Gudina et al., 2021]. Together with the observational model bridging compartment trajectories with empirical data they serve as the foundation for our study. Moreover, we outline Bayesian inference for stochastic differential equation-based models, and explain the chosen techniques of CNF and PF together with the implementation of our comparison workflow. In the second part of the work, we outline the results from comparing the two inference methods on synthetically generated data for the chosen compartmental models and additionally using real-data for the two-variant SEIR model. For the two smaller compartmental models, SIS and SIR, we compared the estimated posterior distributions with a reference posterior obtained by Hamiltonian Monte Carlo (HMC) sampling on a discretized version of the stochastic model. Both methods provide robust and reliable inference results on the stochastic versions of the SIS and SIR model with synthetically generated data, validating their implementation at hand. Analyzing the agreement of the estimated marginal posteriors and model fits based on different synthetic datasets for the two-variant SEIR model, including (partially) missing data, shows that both methods yield good fits to the data at hand. However, ill-conditioning of the model leads to differences in the shape of the marginal posteriors due to difference in parameter space exploration by the inference methods. Applying both methods to an reparametrized version with parameter dimension reduction can overcome these issues and improves posterior alignment. Inference based on the data from Gudina et al. [2021] shows the applicability of both methods to real-world data. A discussion of the results, their limitations and further research directions follows in Section 6.
@@ 2 Stochastic Compartmental Models
Compartmental models are a mathematical framework to describe how individuals in a population move between different states and are fundamental for modeling infectious diseases [Brauer, 2008]. Such a model partitions a population into $d_{X}$ mutually exclusive compartments and is characterized by the vector of compartment sizes $X\in\mathbb{R}^{d_{X}}$ together with the transitions of individuals between these compartments.
Let $\mathcal{T}$ denote the set of admissible transitions. Each transition $(i,j)\in\mathcal{T}$ is characterized by a stoichiometric vector $s_{i,j}\in\mathbb{Z}^{d_{X}}$, which specifies the change in the compartment counts when the transition occurs. For the movement of one individual from compartment $i$ to compartment $j$, $s_{i,j}=e_{j}-e_{i}$, where $e_{i}$ denotes the $i$-th unit vector. The transition intensity is described by a propensity function $a_{i,j}^{N}:\mathcal{S}_{N}\to\mathbb{R}_{+}$, defined through
$\Pr\bigl(X^{N}(t+h)=x+s_{i,j}\mid X^{N}(t)=x\bigr)=a_{i,j}^{N}(x)\,h+o(h),\qquad h\downarrow 0.$
For infection-driven transitions, a common choice is a mass-action propensity of the form
$a_{i,j}^{N}(X)=c_{i,j}X_{i}\sum_{k\in I_{i,j}}\frac{X_{k}}{N},$
where $c_{i,j}>0$ is a rate parameter and $I_{i,j}$ contains the compartments contributing to the force of infection. Other transitions, such as recovery, are typically modeled by linear propensities, for example $a_{i,j}^{N}(x)=c_{i,j}x_{i}$.
The resulting stochastic compartmental model is a continuous-time Markov jump process. In the large-population regime, it can be approximated by the multivariate stochastic differential equation (SDE) [Dragatz, 2006]:
@@
$=\sum_{(i,j)\in\mathcal{T}}a_{i,j}(X(t))s_{i,j}dt+\sum_{(i,j)\in\mathcal{T}}\sqrt{a_{i,j}(X(t))}s_{i,j}s_{i,j}^{T}d\mathbf{B}(t)$
where $\mathbf{B}$ denotes a vector of independent Brownian motions. The representation of compartmental models in terms of multivariate SDEs can capture the randomness of disease spread in their diffusion part $\sigma(X(t))$ and retain a drift component $\mu(X(t))$ for representing the deterministic trend of the population [Andersson and Britton, 2001; Dragatz, 2010]. This makes them a favorable choice for modeling disease outbreaks and dynamics.
In this work, we consider two examples of SDE-based compartmental models, a simple standard Susceptible-Infected-Recovered (SIR) model [Kermack et al., 1927] and an extended two-variant Susceptible-Exposed-Infected-Recovered (SEIR) model [Hethcote, 2000] considering two variants of a single pathogen.
Figure 1: Schematic representations of the models and methods. (a) Graph of compartments and possible transitions with corresponding rate parameters for the Susceptible - Infected - Susceptible (SIS) model. (b) Graph of compartments and possible transitions with corresponding rate parameters for the Susceptible - Infected - Recovered (SIR) model. (c) Graph of compartments and possible transitions with corresponding rate parameters for the Susceptible-Exposed-Infected-Recovered (SEIR) model with virus variants wild-type (wt) and variant (var). (d) Workflow for the assessment of the Bayesian inference.
2.1 SIS Model
The simplest compartmental model is the SIS-model [Kermack et al., 1927] consisting of one Susceptible compartment $S$ and one Infectious compartment $I$, where people can transition between them (Figure 1(a)). This model can be used for infectious diseases that do not confer any long-lasting immunity, such as the common cold or influenza. Following simple mass-action incidence individuals get infected with a rate $\beta SI$ and leave the infective class at rate $\gamma I$ (Table 1).
Table 1: Description of the SIS model. Table including all events, such as initialization and transition of individuals, with their respective rate and model parameters.
Event Description Rate Parameter
Recovery $\mathrm{I}_{\text{}}\rightarrow\mathrm{S}_{\text{}}$ $\gamma\mathrm{I}(t)$ $\gamma^{-1}$
@@ 2.2 SIR Model
The standard SIR model (Figure 1(b)) considers three compartments: $S$ (Susceptible), $I$ (Infectious) and $R$ (Recovered). The model describes the number of individuals in these compartments given the transition parameters $\beta$ (the transmission rate) and $\gamma$ (the recovery rate). Due to its simplicity, the solution of the deterministic SIR model and its properties can be analysed analytically [Harko et al., 2014] and it often serves as an initial model for the modeling of unknown diseases [Daley and Gani, 1999]. For consistency across models, in the following we compare recovery in terms of the mean infectious period, $\gamma^{-1}$, rather than $\gamma$. The events dynamics are described by three compartments and the transitions between them (Table 2).
@@ 2.3 Two-Variant SEIR Model
Many extensions of the SIR model have been developed to account for the heterogeneity of various infectious diseases [Keeling and Rohani, 2008], for example the Susceptible-Exposed-Infectious-Recovered model, in which the Exposed compartment covers the latent period between being infected and becoming infectious. Recently, an extension to the ODE-based SEIR model was introduced to capture the early COVID-19 dynamics in Ethiopia [Gudina et al., 2021]. In addition to a classical SEIR model, this model distinguishes between two virus variants, namely the wild-type and a novel variant, leading to a total of 10 compartments (Figure 1(c)).
The novel variant is assumed to have a 35% longer infectious period, following the original formulation of the two‑variant SEIR model in Gudina et al. [2021]. We adopt this fixed factor to remain consistent with the published model on which our comparison is based. Additionally, immunity from the wild-type variant does not protect against reinfection with the novel variant. However, an infection with the novel variant type confers immunity against subsequent infections with both the wild-type and the novel variant. The dynamics of the compartment sizes are described by initialization, variant entering and transitions between the compartments (Table 3).
The model exhibits non-identifiability issues in the sense that different parameter combinations, lying along certain curves, can generate nearly indistinguishable epidemic trajectories (see Supplementary Information S1.3.6.2). Therefore, in addition to this model, we also consider a reparametrized version of the model with a reduced number of model parameters to be estimated. For this, we fix $\kappa^{-1}=5$, since $\kappa$ seems to participate in the non-identifiable parameter combinations and use the following redefined parameters
$r_{0}=\beta\cdot\gamma^{-1},\quad e_{0}=\kappa^{-1}+\gamma^{-1},\quad s_{0}=s\cdot\gamma^{-1},$ (1)